2n2-31+1 b. lim cos 1. Find the following limits: a. lim 1 00 --- 3n2+4 n-00
Answer #1

1 a lim 2 n²-3n+1 bin nyo 3n2tu. nos a (2-3 44 + 1/ 42) (3th ) 3to lim Cos ñ = t-20 no het t= -2 a n-icoito tim Cost = Coro =